NottheMonga2087:

The volume, V, of a rectangular prism is determined using the formula V=lwh, where l is the length, w is the width, and h is the height of the prism. Using this formula, what is the width of a rectangular prism that has a volume of 138.24 cubic inches, a height of 9.6 inches, and a length of 3.2 inches? Round to the nearest tenth. 0.2 inches 4.5 inches 46.1 inches 414.7 inches

sophia333:

volume=l×b×h 138.24=3.2×b×9.6 138.24=30.72×b b=138.72/30.72 b=4.5 hence breadth = 4.5 inches

NottheMonga2087:

thank you!!

sophia333:

Np :)

sophia333:

If ya wouldnt mind could i has a medal :>